18. The position of a particle moving along the x-axis is given by x-30t-30t, where x is achieves its maximum speed in the positive x-direction? A) 42 m B) 9.8 m C) 3 m D) 27 m E) 5.6 m
Answer #1

given

x = 30t2 - 30 t4

we know that

v = dx/dt = d(30t2 - 30 t4)/dt

v = 60 t - 120 t3

for maximum velocity dv/dt = 0

d(60 t - 120 t3)/dt = 0

60 - 360 t2 = 0

t2 = 60/360 = 1/6

t = 1/ sec

so this instant X = ?

put t = 1/ in x

x =  30t2 - 30 t4

= 30 (1/)2 - 30(1/)4

x = 4.2 m

answer